Frank borrows 10000 dollars at 8.2 percent annual interest, compounded once per year, when he begins college. Four years later, how much will he owe? (Assume that he makes no payments during the four years.)
Answer by mananth(16946)

Principal P = 10000
Amount= A
years=n 4.00
compounded 1 times a year t
Rate = 8.20 0.082
Amount = P*((n+r)/n)^n*t

Amount = = 10000 *(1 +0.08)^ 4*1
Amount =10000 *(1+0.082)^4
10000 *( 1.08 )^ 4
Amount = 13705.95